### v1

# Absorption and scattering of massless scalar waves by Frolov black holes

Abstract
We comprehensively investigate the absorption and scattering of massless scalar waves by Frolov black holes, which is a class of regularization of the Reissner–Nordström spacetime. By analyzing the null geodesics, we determine the photon sphere radius and the critical impact parameter, deriving the geometric capture cross section and the classical differential scattering cross section. Utilizing the partial-wave method, we numerically compute the absorption and scattering cross sections across a broad frequency range. Our numerical results show excellent agreement with the low-frequency limit and the high-frequency sinc approximation, as well as with the semiclassical glory approximation. We analyze the dependence of the spectra on the charge and the regularization parameter (Hubble length). Under the horizon-radius normalization, we observe that the total absorption cross section increases with the regularization parameter, a behavior we attribute to the variation of the dimensionless mass. Furthermore, by comparing Frolov, Reissner–Nordström, and Hayward black holes, we demonstrate that their absorption and scattering patterns are nearly indistinguishable when their critical impact parameters or glory impact parameters are matched. This indicates that the photon sphere geometry predominantly governs scalar field interactions, while the detailed core structure plays a secondary role. Our work underscores the potential of wave-based probes to test regular black hole geometries and their observable imprints.
Regular black holes, Massless scalar wave, Scattering, Absorption
I Introduction
With the advent of gravitational-wave astronomy and black hole imaging[40, 39, 12, 47, 29, 1], understanding the interaction between waves and black holes has become more crucial than ever. The scattering and absorption of fields by black holes provide powerful probes of spacetime structure and tests of general relativity[2, 44]. Related wave-scattering phenomena have also been discussed in alternative gravitational settings, e.g., for (quasi-)spherical gravitational waves scattered from black strings in massive gravity[52]. Propagating fields leave characteristic imprints in their absorption and differential scattering cross sections, which depend both on universal near-horizon physics and on the detailed geometry of the photon sphere and its unstable null orbits. At low frequencies these observables typically display universal behaviour, largely insensitive to the fine details of the spacetime [14], whereas at high frequencies they are controlled by the capture properties of null geodesics and by the structure of the photon sphere, which also underlies the phenomenology of black hole shadows and photon rings as observed in M87* and Sgr A* [19, 20]. Wave-optics diffraction phenomena can also arise in black-hole scattering, such as the Poisson–Arago spot discussed for gravitational waves[51].
For a minimally coupled massless scalar field, this picture is well understood in benchmark geometries. In the zero-frequency limit the total absorption cross section approaches the area of the event horizon [24, 25], while in the high-frequency regime it oscillates around the geometric capture cross section, with an oscillatory pattern that can be related to the properties of the unstable photon orbit and the associated Regge poles of the scattering matrix [17, 46]. These features have been established in detail for Schwarzschild and Reissner–Nordström black holes through a combination of partial-wave analysis and explicit geodesic control [22, 13]. Full-wave computations of scalar scattering in nontrivial charged backgrounds, together with comparisons to geodesic and Regge-pole/glory approximations, have been carried out for charged dilatonic black holes [28].
The quest to resolve the central singularities predicted by general relativity has motivated the study of regular black holes (RBHs) [3, 31, 9, 7]. Many such solutions arise in nonlinear electrodynamics and modify the inner structure while preserving asymptotic flatness [4, 5, 6]. In particular, the Bardeen and Hayward spacetimes [8, 23] provide simple models in which regular cores and effective charges reshape the effective potential and photon-sphere data, leading to characteristic changes in greybody factors and large-angle scattering patterns [37, 38].
As a charged generalization of the Hayward solution, the Frolov black hole [21] has recently attracted considerable interest. Its thermodynamic properties, quasinormal modes and shadows have been investigated in a variety of contexts [43, 48, 30]. However, to the best of our knowledge, a systematic study of massless-scalar absorption and scattering in this spacetime, covering analytic low- and high-frequency limits and full-frequency numerical results, is still lacking.
Against this backdrop, the present work develops a comprehensive analysis of massless scalar absorption and scattering by Frolov black holes. At the classical level we study null geodesics, identifying the photon sphere and the associated critical impact parameters, and we derive the weak-deflection and backward glory approximations for the differential cross section. Within the partial-wave formalism we obtain the radial equation and its effective potential, and we derive the usual low-frequency limit and a high-frequency sinc approximation expressed in terms of the critical impact parameter and Lyapunov exponent. We then compute the absorption and differential scattering cross sections numerically over a broad frequency range, comparing them with the analytic approximations and exploring how the two Frolov parameters jointly affect the spectra. In particular, we examine cases in which different parameter choices yield very similar absorption and scattering patterns, and relate this behaviour to the structure of the effective potential and the underlying geodesics.
The remainder of this paper is organized as follows. In Sec. II we review the Frolov spacetime, analyse null geodesics and discuss the corresponding classical capture and scattering properties. In Sec. III we present the partial-wave formalism for a massless scalar field, derive the effective potential, and obtain the analytic low- and high-frequency limits for the absorption and scattering cross sections. Sec. IV is devoted to the numerical results: we summarize our normalization choices, describe the numerical implementation, and present the full-frequency absorption and scattering spectra for a range of Frolov parameters, including comparisons with Reissner–Nordström and Hayward black holes at matched impact parameters. We conclude our results in Sec. V. Throughout this work we use natural units with $G=c=\hbar=1$ and metric signature (,+,+,+)(-,+,+,+). For analytic expressions we keep the mass parameter $M$ explicit to facilitate comparison with the existing literature, whereas in our numerical calculations and plots we adopt the horizon-radius normalization $r_{h}=1$, defined by $f(r_{h})=0$.
II Classical Analysis
We consider the Frolov BH, which is a static regular black hole, whose line element is given by [21]
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (1)
where
$f(r)=1-\frac{r^{2}(2Mr-Q^{2})}{r^{4}+(2Mr+Q^{2})\alpha^{2}}.$ (2)
Here $M$ and $Q$ respectively denote the mass and charge parameter of the black hole. The regularized parameter $\alpha$ represents the Hubble length, which is related to the effective cosmological constant $\Lambda_{\text{eff}}=3/\alpha^{2}$, characterize the curvature core of the Frolov BH. For keeping the event horizon surviving, the Hubble length is bounded with $\alpha\leq\sqrt{16/27}M$.
As $r\to\infty$, the spacetime is asymptotically Reissner–Nordström $f(r)\approx f_{\text{RN}}(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}$, where near the center $r\to 0$ the expansion $f(r)=1-\frac{\Lambda_{\text{eff}}}{3}r^{2}+\cdots$ signals regularity.
II.1 Null geodesics and geometrical limits
The classical Lagrangian related to the null geodesic trajectory is given by
$\mathcal{L}=\frac{1}{2}g_{\mu\nu}\frac{dx^{\mu}}{d\lambda}\frac{dx^{\nu}}{d\lambda}=0.$ (3)
Let us constrain the motion of particles on the equatorial plane without loss of generality, the equation above gives
$f(r)\dot{t}^{2}-\frac{\dot{r}^{2}}{f(r)}-r^{2}\dot{\phi}^{2}=0.$ (4)
It’s clear that there exist two conserved quantities
$E=\frac{\partial\mathcal{L}}{\partial\dot{t}}=-f\dot{t}\quad,\quad L=-\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=r^{2}\dot{\phi},$ (5)
where $E$ and $L$ are respectively the energy and the angular momentum of the massless particle. Then the geodesic equation can be written as
$\dot{r}^{2}=f^{2}\dot{t}^{2}-fr\dot{\phi}^{2}=E^{2}-f\frac{L^{2}}{r^{2}},$ (6)
here we can write the equation of motion Eq.(6) in the form
$\dot{r}^{2}+V_{\text{eff}}=E^{2}$ (7)
by defining
$V_{\text{eff}}=f(r)\frac{L^{2}}{r^{2}}=\left[1-\frac{r^{2}(2Mr-Q^{2})}{r^{4}+(2Mr+Q^{2})\alpha^{2}}\right]\frac{L^{2}}{r^{2}}$ (8)
In Fig.2, we display the effective potential for massless particles in Frolov BHs with different parameters for illustration.
After using the impact parameter $b\equiv L/E$ as the ratio of angular momentum and energy, we can obtain the orbit equation
$\left(\frac{dr}{d\phi}\right)^{2}=\frac{\dot{r}^{2}}{\dot{\phi}^{2}}=r^{4}\left(\frac{1}{b^{2}}-\frac{f}{r^{2}}\right).$ (9)
It is convenient to introduce $u\equiv 1/r$, the orbit equation can be written as
$\left(\frac{du}{d\phi}\right)^{2}=\left(\frac{du}{dr}\right)^{2}\left(\frac{dr}{d\phi}\right)^{2}=\mathcal{U}(u)=\frac{1}{b^{2}}-u^{2}f,$ (10)
where $f=f(1/u)$. By differentiating Eq.(10), we get
$\frac{d^{2}u}{d\phi^{2}}=-\frac{u^{2}}{2}\frac{df}{du}-uf$ (11)
For the Frolov spacetime,
$f(r)=1-\frac{(2Mr-Q^{2})r^{2}}{r^{4}+(2Mr+Q^{2})\alpha^{2}},$ (12)
two orbit equations read
$\left(\frac{du}{d\phi}\right)^{2}=\frac{1}{b^{2}}-u^{2}+\frac{u^{3}\left(2M-Q^{2}u\right)}{\alpha^{2}u^{3}\left(2M+Q^{2}u\right)+1},$ (13)
$\frac{d^{2}u}{d\phi^{2}}\;=\;-\,\frac{P(u)}{B(u)^{2}},$ (14)
where
$B(u)=1+\alpha^{2}u^{3}\bigl(2M+Q^{2}u\bigr),$ (15)
$P(u)={}$ $u\Bigl[1-3Mu+2Q^{2}u^{2}+4\alpha^{2}Mu^{3}+2\alpha^{2}Q^{2}u^{4}+2\alpha^{2}MQ^{2}u^{5}$ (16)
$\hskip 55.00008pt+\,\alpha^{4}\bigl(4M^{2}u^{6}+4MQ^{2}u^{7}+Q^{4}u^{8}\bigr)\Bigr].$
It is obvious that, for the curvature-free case ($\alpha=0$),
$f=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}=1-2Mu+Q^{2}u^{2},$ (17)
the equations are consistent with previous work on RN BHs [13, 11]
$\left(\frac{du}{d\phi}\right)^{2}=\frac{1}{b^{2}}-u^{2}+2Mu^{3}-Q^{2}u^{4},$ (18)
$\frac{d^{2}u}{d\phi^{2}}=-u+3Mu^{2}-2Q^{2}u^{3}.$ (19)
From the orbit equation Eq.(10), one sees $\mathcal{U}(u)\geq 0$ along the orbit. There is a critical value of the impact parameter $b_{c}$, which related to the critical radius $r_{c}$ of the unstable circular orbit, namely
$b_{c}=\frac{L_{c}}{E_{c}}=\frac{r_{c}}{\sqrt{f(r_{c})}},\quad 2f(r_{c})-r_{c}\frac{df(r_{c})}{dr}=0.$ (20)
The trajectories can be obtained by numerically solving the orbit equation (9). In Fig.1, we give the null geodesic orbits in Frolov spacetime with several parameters. In particular, the photon-sphere data entering $b_{c}$ also control the CAM ingredients, like Regge frequency and damping, as explicitly demonstrated in Ref. [34].
These null-geodesic quantities, in particular the photon-sphere radius and the critical impact parameter $b_{c}$, also underlie the optical appearance of compact objects. Recent analyses of photon dynamics and shadows in rotating spacetimes with parity-odd hair provide closely related geodesic diagnostics [26]. See also Ref. [42] for a systematic discussion of null-particle dynamics and shadows in a general rotating black-hole geometry. Light rings and lensing diagnostics are not restricted to black holes; for related discussions in rotating boson-star spacetimes with parity-odd features, see Ref. [27].
Equation (20) partitions incoming null geodesics into two classes: for $|b|<b_{c}$ the trajectory crosses the potential barrier and is captured, whereas for $|b|>b_{c}$ it turns back to infinity. The geometric cross section, which is the classical capture cross section of geodesics, is well known[49]
$\sigma_{\rm geo}=\pi b_{c}^{2}.$ (21)
In high frequency limit, the absorption cross section can be written as the sinc approximation, which involves leading term $\sigma_{\rm geo}$ and the oscillatory correction[46, 17]
$\sigma_{\rm abs}^{\rm hf}\approx\sigma_{\rm geo}\left[1-8\pi\frac{\Lambda_{c}}{\Omega_{c}}e^{-\pi\Lambda_{c}/\Omega_{c}}\operatorname{sinc}(2\pi b_{c}\omega)\right],$ (22)
where $\operatorname{sinc}(x)\equiv\sin(x)/x$. And the photon-sphere angular frequency and Lyapunov exponent[10], $\Omega_{c}=1/b_{c}$ is the angular frequency of the unstable photon circular orbit at $r_{c}$. Lyapunov exponent of these circular null geodesics is given by[10]
$\Omega_{c}=\frac{1}{b_{c}}=\frac{\sqrt{f(r_{c})}}{r_{c}},\quad\Lambda_{c}=\sqrt{\left.\frac{L^{2}}{2\dot{t}^{2}}V^{\prime\prime}(r)\right|_{r_{c}}}=\sqrt{\frac{f(r_{c})\Bigl[2f(r_{c})-r_{c}^{2}f^{\prime\prime}(r_{c})\Bigr]}{2r_{c}^{2}}},$ (23)
will anchor the oscillatory pattern of the high-frequency absorption in Sec. IV.2.
(a) Null geodesics for different $Q/r_{h}$ with $\alpha/r_{h}=0.3$
(b) Null geodesics for different $\alpha/r_{h}$ with $Q/r_{h}=0.3$
Figure 1: Null geodesics for massless particles in the background of Frolov BHs with different values of $Q$ and $\alpha$ for a fixed $b=3r_{h}$. The trajectories can be derived numerically by orbit equation Eq. (10). We have set the initial condition $r_{\rm inf}$ at $100r_{h}$.
II.2 Weak deflection and Classical scattering
Having established the geometrical-optics limit via the photon sphere ($r_{c},b_{c}$) and the capture cross section $\sigma_{\rm geo}$ in the previous subsection, we now focus on the rays that return to infinity, i.e. the scattering branch $b>b_{c}$. In the scattering case, the deflection angle can be obtained by integrating Eq.(10)
$\Theta(b)=2\int_{0}^{u_{0}}\frac{du}{\sqrt{\mathcal{U}(u)}}-\pi=2\int_{0}^{u_{0}}du\left(\frac{1}{b^{2}}-u^{2}f\right)^{-1/2}-\pi.$ (24)
In the weak-field limit, the deflection angle can be expanded in power of $b^{-1}$
$\Theta\approx\frac{4M}{b}+\frac{3\pi}{4b^{2}}\left(5M^{2}-Q^{2}\right)+\mathcal{O}\left(\frac{1}{b}\right)^{3},$ (25)
which coincides with the RN result. Expanding the lapse at large radii yields
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\alpha^{2}\!\left(\frac{4M^{2}}{r^{4}}-\frac{Q^{4}}{r^{6}}\right)+\cdots,$ (26)
so the $\alpha$-dependent corrections start at $O(r^{-4})$ and therefore enter the weak-deflection series only at $O(b^{-4})$. This is consistent with the weak-deflection expansion obtained via the Gauss-Bonnet approach in Ref. [30] and with the Hayward case [16].
The classical differential cross section is given by [44]
$\left.\frac{d\sigma_{\rm sc}}{d\Omega}\right|_{\rm cl}=\frac{b}{\sin\theta}\left|\frac{db}{d\theta}\right|.$ (27)
From Eq.(25) and (27), we conclude that the classical differential scattering cross section for small scattering angles is
$\left.\frac{d\sigma_{\rm sc}}{d\Omega}\right|_{\rm cl}=\frac{16M^{2}}{\Theta^{4}}+\frac{15\pi M^{2}}{4\Theta^{3}}-\frac{3\pi Q^{2}}{4\Theta^{3}}+\mathcal{O}\left(\frac{1}{\Theta}\right)^{2}.$ (28)
It is obvious that the Hubble length $\alpha$ has no effect on the dominant term of classical differential scattering cross section.
For the glory approximation of the scalar scattering cross sections, the backward glory is defined by $\Theta(b_{g})=\pi$, in the spacetime of a static and symmetric BH, the glory approximation of the scalar scattering cross sections is [41]
$\left.\frac{d\sigma_{sc}}{d\Omega}\right|_{\theta\approx\pi}\approx 2\pi\omega b_{g}^{2}\left|\frac{db}{d\theta}\right|_{\theta=\pi}\left[J_{0}\left(\omega b_{g}\sin\theta\right)\right]^{2},$ (29)
where $\omega$ is the frequency of the scalar wave, $b_{g}$ is the impact parameter that corresponds to a deflection angle of $\pi$, $J_{0}(x)$ is a Bessel function of the first kind. Imposing the backward-scattering condition $\theta=\pi+2n\pi$ ($n=0,1,2,\ldots$) yields a discrete set of glory impact parameters $b_{g}^{n}$, where $n$ counts the number of windings of the null geodesic around the photon sphere. In practice, the differential cross section near $\theta\simeq\pi$ is dominated by the $n=0$ branch, while higher-winding contributions are strongly suppressed [38, 15]. Although the semi-classical glory approximation (29) is formally derived for the high-frequency regime, it remains in very good agreement with the numerical results down to intermediate frequencies [13].
III Partial-Wave Method
III.1 Massless scalar field
First we consider a minimally coupled massless scalar obeying the Klein-Gordon equation
$\nabla_{\mu}\nabla^{\mu}\Phi=\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Phi\right)=0.$ (30)
With the standard separation
$\Phi=\frac{\psi_{\omega l}(r)}{r}Y_{lm}(\theta,\phi)e^{-i\omega t},$ (31)
the radial function satisfies a RW-type equation in the tortoise coordinate
$\frac{d^{2}}{dr_{*}^{2}}\psi_{\omega l}+\left[\omega^{2}-V_{\rm eff}(r)\right]\psi_{\omega l}=0,$ (32)
with effective potential
$V_{\rm eff}(r)=f(r)\left[\frac{f^{\prime}(r)}{r}+\frac{l(l+1)}{r^{2}}\right],$ (33)
where tortoise coordinate $r_{*}$ is defined as $f(r)dr_{*}=dr$.
Figure 2: The effective potential of a Frolov BH with the parameters $Q/r_{h}=0.3$ and $\alpha/r_{h}=0.3$.
Figure 3: Effective potential $V_{\text{eff}}(r)$ for the $l=0$ mode of Frolov BHs.
In Fig. 2 we plot the effective potential for a "mild" Frolov black hole with parameters chosen as $Q/r_{h}=0.3$ and $\alpha/r_{h}=0.3$. Fig. 3 compares the effective potential, with $l=0$ for various parameters. As we can see, no matter which parameter is fixed, the maximum of the potential always decreases as the other parameter increases. This apparent discrepancy with the trend reported in works that adopt the $M=1$ normalization for RN and Hayward black holes arises solely from the different choice of units; a detailed discussion of the relation between the $r_{h}=1$ and $M=1$ conventions is given in Sec. IV.
As $r_{*}\to\pm\infty$ the effective potential vanishes, $V_{\rm eff}\to 0$, so Eq. (32) reduces to $\psi^{\prime\prime}+\omega^{2}\psi=0$ and the solutions are plane waves. It is convenient to express the large-argument behavior with the spherical Hankel functions $h_{l}^{(1)}$ and $h_{l}^{(2)}$, which satisfy $h_{l}^{(1)}(z)\sim(-i)^{\,l+1}e^{iz}/z$ and $h_{l}^{(2)}(z)\sim i^{\,l+1}e^{-iz}/z$ as $z\to\infty$. Imposing purely ingoing behavior at the horizon and an incident-plus-outgoing superposition at infinity yields the boundary conditions
$\psi_{\omega l}\sim\begin{cases}A^{\rm tr}_{\omega l}\,e^{-i\omega r_{*}},&r_{*}\to-\infty,\\[2.0pt] A^{\rm in}_{\omega l}\,e^{-i\omega r_{*}}+A^{\rm out}_{\omega l}\,e^{+i\omega r_{*}},&r_{*}\to+\infty,\end{cases}$ (34)
and flux conservation gives $|A^{\rm tr}_{\omega l}|^{2}+|A^{\rm out}_{\omega l}|^{2}=|A^{\rm in}_{\omega l}|^{2}$.
It is convenient to introduce the partial-wave $S$-matrix, defined by the ratio of the outgoing and ingoing amplitudes
$S_{l}(\omega)=(-1)^{l+1}\frac{A^{\rm out}_{\omega l}}{A^{\rm in}_{\omega l}}.$ (35)
The absorption probability of the $l$-th mode is given by
$\Gamma_{l}(\omega)=1-|S_{l}|^{2}=\frac{|A^{\rm tr}_{\omega l}|^{2}}{|A^{\rm in}_{\omega l}|^{2}},$ (36)
which characterizes the transmission of the scalar wave through the effective potential barrier.
III.2 Absorption cross sections
The total absorption cross section is the sum over partial contributions
$\sigma_{\rm abs}=\sum_{l=0}^{\infty}\sigma_{\rm abs}^{(l)},$ (37)
with
$\sigma_{\rm abs}^{(l)}=\frac{\pi}{\omega^{2}}(2l+1)\Gamma_{l}=\frac{\pi}{\omega^{2}}(2l+1)\left(1-\left|\frac{A_{\omega l}^{\rm out}}{A_{\omega l}^{\rm in}}\right|^{2}\right).$ (38)
In the low-frequency limit only the s-wave survives, then the cross section will approach to the area of BH horizon, this universality first established in Ref. [14] and extended to stationary geometries in Ref. [24]. But in high-frequency regime, capture dominates and the cross section rapidly oscillating around the geometric cross section $\sigma_{\rm geo}$, which we have discussed in Sec. II.1. Here we give the expression of absorption cross section in two asymptotic limits
$\sigma_{\rm abs}^{\rm lf}\approx A_{H},\quad\sigma_{\rm abs}^{\rm hf}\approx\pi b_{c}^{2}-8\pi^{2}\frac{\Lambda_{c}}{\Omega_{c}^{3}}e^{-\pi\Lambda_{c}/\Omega_{c}}\operatorname{sinc}(2\pi b_{c}\omega).$ (39)
III.3 Scattering cross section
The scattering amplitude is given by
$g(\theta)=\frac{1}{2i\omega}\sum_{l=0}^{\infty}(2l+1)\left[e^{2i\delta_{l}}-1\right]P_{l}(\cos\theta)$ (40)
and the differential scattering cross section is
$\frac{d\sigma_{\rm sc}}{d\Omega}=\bigl|g(\theta)\bigr|^{2},$ (41)
in this description the phase shifts defined in Eq. (35) encode the same information as the classical deflection function discussed in Sec. II.
Our numerical evaluation of $S_{l}(\omega)$ and the resulting $\sigma_{\rm abs}(\omega)$ and $d\sigma_{\rm sc}/d\Omega$, together with comparisons to the low/high-frequency limits and the classical benchmarks of Sec. II, are presented in the next section.
IV Numerical Results
In this section, we present a series of numerical results concerning the absorption and scattering cross sections of massless scalar field in the background of Frolov spacetimes with variable parameters.
IV.1 Numerical setup
Given the complexity of Frolov black holes, directly solving the horizon by setting the metric function (2) equal to zero would introduce additional computational complexity and numerical uncertainty. For this reason, unlike previous studies on absorption and scattering of BHs, we adopt the normalization by setting $r_{h}=1$ instead of working with the commonly used $M=1$. It is convenient to introduce the normalized parameter
$q\equiv\frac{Q}{r_{h}},\qquad\ell\equiv\frac{\alpha}{r_{h}},\qquad m\equiv\frac{M}{r_{h}},$ (42)
in place of the original $Q$, $\alpha$ and $M$.
This choice imposes nontrivial constraints on the admissible values of the black hole parameters. The first constraint requires that the charge $q$ satisfy the following inequality with the Hubble length $\ell$:
$q^{2}-\ell^{4}q^{2}+4\ell^{2}q^{2}+3\ell^{2}\leq 1.$ (43)
In Fig. 4, we show the parameter space in the $(\ell^{2},q^{2})$ plane. The background color encodes the function $E(\ell^{2},q^{2})=-1+q^{2}-\ell^{4}q^{2}+\ell^{2}(3+4q^{2})$, with the blue region $E\leq 0$ corresponding to configurations that admit an event horizon, while the complementary region $E\geq 0$ represents horizonless geometries. The solid black curve marks the locus $E=0$, where the two horizons merge and the Frolov BH becomes extremal; its intercepts at $(\ell^{2}=0,q^{2}=1)$ and $(\ell^{2}=1/3,q^{2}=0)$ reproduce, respectively, the extremal Reissner–Nordström and extremal Hayward limits. It is worth noting that, since we employ the horizon radius as our fundamental length scale, the limiting value of the regularizing parameter in the uncharged case is now $\ell=\sqrt{1/3}$, rather than the more familiar $\sqrt{16/27}$ quoted in the $M=1$ convention.
Figure 4: Parameter space of Frolov BHs in the $(\ell^{2},q^{2})$ plane. The blue region corresponds to configurations with an event horizon, and the solid black curve marks the extremal limit.
Within this domain, the condition $f(r_{h}=1)=0$ fixes the dimensionless mass as a function of the remaining parameters
$m(q,\ell)=\frac{1+(1+\ell^{2})q^{2}}{2(1-\ell^{2})}.$ (44)
Figure 5 displays the behaviour of $m$ as a function of $q^{2}$ and $\ell^{2}$ within the region of parameter space that admits an event horizon, which we discussed in the last paragraph. One can see that $m$ increases monotonically with both $q$ and $\ell$, so different choices of $(q,\ell)$ along the allowed domain correspond to Frolov BHs of different mass, even though the horizon radius is kept fixed. As a consequence, varying $(q,\ell)$ at fixed $r_{h}=1$ is not directly equivalent to varying the charge $Q$ or the Hubble length $\alpha$ at fixed mass in the usual $M=1$ normalization. This observation will be important when interpreting, in the next subsections, the dependence of the absorption and scattering cross sections on $(q,\ell)$ and when comparing our results with previous absorption scattering research in references.
Figure 5: Normalized mass $m$ of Frolov BHs for a fixed horizon radius $r_{h}=1$, restricted to the region admitting an event horizon.
Finally we numerically solve Eq. (32) from the point very close to the event horizon to a region far from the BHs, then match the numerical solutions with the boundary condition given in Eq. (34) to compute the coefficients. In calculations of absorption and scattering cross sections, we perform sums on angular momentum of the scalar wave. For the absorption case, we set $l=10$, and consider $l=50$ for the scattering case. To optimize the convergence of the differential scattering cross section at small scattering angles, we employed the well-known YRW method[50, 18]. Related convergence issues of the partial-wave expansion near the optical axis, and a complementary resolution based on finite-radius wavefunctions (avoiding the problematic asymptotic expansion), were analyzed in a rigorous manner in Ref. [36].
IV.2 Scalar absorption
In Fig. 6, we plot the total absorption cross sections (ACS) of massless scalar waves in the Frolov spacetime. The left panel shows the dependence on the Hubble length $\ell$ for fixed charge $Q$, whereas the right panel displays the dependence on $Q$ at fixed $\ell$, illustrating how the ACS changes as the parameters are varied toward the extremal regime.
Figure 6: Total ACS of massless scalar fields in Frolov spacetime. The numerical results (solid lines) are compared with the high-frequency sinc approximation (dotted lines) and the geometric capture limit (dashed lines). Panels display the dependence on the Hubble length $\ell$ at fixed charge $q=0.5$ (left) and on $q$ at fixed $\ell=0.3$ (right)
As we can see in this figure, the numerical results exhibit the expected asymptotic behaviors at both frequency ends. In the low-frequency limit ($\omega r\to 0$), the cross sections universally converge to the horizon area $\sigma_{\rm abs}\to 4\pi r_{h}^{2}$, while in the high-frequency regime, the spectra oscillate around the geometric cross section $\sigma_{\rm abs}\to\sigma_{\rm geo}=\pi b_{c}^{2}$ with a pattern that is well-described by the semiclassical sinc approximation. Apart from these limits, we observe that the overall amplitude of the ACS typically increases as we increase either $q$ or $\ell$. This trend appears opposite to what is reported for RN and Hayward black holes in the standard $M=1$ normalization, where the total ACS usually decreases as the charge or regularizing parameter grows [13, 16]. However, in our $r_{h}=1$ convention the dimensionless mass $m(q,\ell)$ also increases across the allowed parameter space (see Sec. IV and Fig. 5), so part of the apparent growth of the ACS is simply reflecting the change of the overall mass scale rather than a genuine reversal of the underlying geometric effect.
Then we compare the partial ACSs of scalar wave in Frolov spacetime with the same parameters in Fig. 7 as a complementary view of Fig. 6. These plots show explicitly that the parameter dependence of the total ACS is mainly driven by the lowest modes, in particular the $s$-wave ($l=0$), while higher-$l$ modes provide comparatively smaller influence.
Figure 7: Partial absorption cross sections of massless scalar waves in the Frolov spacetime for the lowest angular-momentum modes $l=0,1,2$. The left panel shows the dependence on the Hubble length $\ell$ for fixed charge $Q=0.5$, whereas the right panel displays the dependence on $Q$ with fixed $\ell=0.3$.
IV.3 Scattering cross section
In this subsection we discuss the scattering of massless scalar waves by Frolov BHs. The differential scattering cross section is computed from the partial-wave series in Eq. (41), with phase shifts obtained from the numerical integration of the radial equation (32).
First we compare our numerical results for the differential scattering cross sections (DCS) of massless scalar waves in the Frolov spacetime with approximations. It is clearly obtained that the DCS oscillating around the classical DCS, which we have derived in Eq. (28). The pattern agrees well with the classical analytical results in small scattering angle, and is well described by glory approximation(Eq. (29)) in the region $\theta\approx\pi$.
Figure 8: Comparison between the numerical and the approximate analytical results for the DCS of massless scalar fields in the Frolov spacetime. Here we consider $q=0.3$ and $\ell=0.3$ as parameters for frequency $\omega r_{h}=4$.
Then Fig. 9 displays the differential scattering cross sections (DCS) for massless scalar waves with frequency $\omega r_{h}=3$ in the background of Frolov spacetime, which show the dependency of DCS on the parameters. The left panel shows the effect of varying $\ell$ at fixed charge, while the right panel shows the effect of varying $q$ at fixed $\ell$, where the values chosen are consistent with we used in the absorption work. One can see that the small-angle cross section is much more sensitive to changes in $q$ than to changes in $\ell$. The behavior is in agreement with the formula in Eq. (28), where the charge enters the subdominating $\Theta^{-3}$ term, whereas the dependence on $\ell$ appears only through higher-orders corrections that is neglected in the approximation. However at large angles, the effect of $\ell$ becomes comparable to that of $q$, which they have a significant effect upon DCS. We also found that the interference fringe width decreases with the increase of two parameters $q$ and $\ell$ for given frequency $\omega r_{h}$. This behavior is opposite to what is typically found for Reissner–Nordström and Hayward black holes when the mass is kept fixed[13, 16]. This difference arises from our choice of units, where the horizon radius rather than mass is used as the fundamental scale.
Figure 9: Differential scattering cross section for Frolov spacetime for different values of $q$ and $\ell$, at a representative frequency $\omega r_{h}=3$.
IV.4 Similar features for different Frolov BHs
The intriguing work of de Paula et al. [16] investigated under which circumstances the absorption and scattering spectra of Hayward BHs and RN BHs can become almost indistinguishable.
In the absorption problem, they let the critical impact parameters coincide by adjusting appropriate parameters, $b_{c}^{\rm H}(\alpha)=b_{c}^{\rm RN}(Q)$, and showed that, under this condition, the corresponding total ACS are very similar over the whole frequency range. In the scattering case, they instead matched the impact parameter of the backscattered null geodesics, requiring $b_{g}^{\rm H}(\alpha)=b_{g}^{\rm RN}(Q)$, and found that the DCS patterns of two black holes are also nearly identical.
Motivated by this observation, we now ask whether a similar correspondence persists when comparing RN, Hayward and Frolov black holes with the same impact parameters. Since the Frolov spacetime reduces to the RN solution and Hayward black hole in appropriate limits, curves of constant impact parameters $b(\ell,q)$ in the $(\ell^{2},q^{2})$ plane [cf. Fig. 4] naturally interpolate between these two limiting geometries. Along a given iso-$b$ curve there thus exists a continuous family of Frolov BHs sharing the same impact parameter.
To assess how far the absorption and scattering spectra are influenced by these impact parameters, we focus on the lines with $b_{c}\simeq 3\,r_{h}$ and $b_{g}\simeq 3\,r_{h}$, and select three representative configurations on each line: an Reissner–Nordström black hole, a Hayward black hole and an intermediate Frolov case. As shown in Fig. 10(a), we plot the ACSs for three cases with $b_{c}\simeq 3\,r_{h}$ over the whole frequency range. Moreover, Fig. 10(b) illustrates the comparison between the DCSs in the background of three different spacetimes.
(a) The total absorption cross section (ACS)
(b) The differential scattering cross section (DCS)
Figure 10: ACSs and DCSs of Reissner–Nordström, Hayward and Frolov black holes with matched impact parameters. Panel (a) shows configurations with critical impact parameter $b_{c}\simeq 3\,r_{h}$, whereas panel (b) refers to configurations with glory impact parameter $b_{g}\simeq 3\,r_{h}$.
We exhibit, in both panels of Fig. 10, that the ACSs and DCSs for three chosen black hole are basically coincide in the whole frequency/angle range. While the agreement in the asymptotic limits is guaranteed by our parameter matching strategy (fixing $r_{h}$ for the low-frequency limit and $b_{c}/b_{g}$ for the high-frequency limit), the nearly perfect overlap in the intermediate regime is non-trivial. This implies that the iso-spectrality is not merely an artifact of the limits, but stems from the fact that the associated effective potentials possess almost identical barrier shapes around the photon sphere. Consequently, the Frolov, Reissner–Nordström, and Hayward black holes become practically indistinguishable via massless scalar scattering in these configurations.
V Conclusion
The absorption and scattering of fields by black holes serve as a powerful probe for exploring the fundamental nature of spacetime geometry and horizon structure. Regular black holes, which smooth out central singularities, have been shown to leave distinct signatures in the scalar absorption cross sections and differential scattering cross sections across several specific models[37, 45]. However, a systematic investigation into how the two key parameters of the Frolov spacetime jointly influence these observable quantities over a broad frequency range has remained lacking prior to this work.
In this work, we first conducted a comprehensive null geodesic analysis to confirm the existence of the event horizon in Frolov black holes, pinpoint the location of the photon sphere, and calculate the critical impact parameters that govern the capture process and glory scattering. On the wave physics front, we employed the partial-wave method to investigate minimally coupled massless scalar fields. By solving the radial equation with appropriately imposed boundary conditions, we derived the partial absorption probabilities, total and partial absorption cross sections, as well as the differential scattering cross section. Across the entire parameter space explored, the numerically computed cross sections smoothly bridge the universal low-frequency regime and the high-frequency geometric optics regime, demonstrating excellent agreement with classical approximations and glory approximations within their respective valid domains.
In the numerical analysis we adopted the normalization $r_{h}=1$ and mapped the allowed region in the $(\ell^{2},q^{2})$ plane where an event horizon exists. In this parametrization the normalized mass $m$ grows monotonically with both $q$ and $\ell$, so changing these parameters at fixed $r_{h}$ does not correspond to the same effect on spacetime as in the more common $M=1$ convention. This explains why, in our results, the total absorption cross section and the height of the effective potential barrier increase as either parameter is raised, even though related studies of Reissner–Nordström and Hayward black holes at fixed mass report a decreasing trend. The partial–wave decomposition further shows that the lowest mode dominate the low–frequency absorption and drive most of the parameter dependence, while variations of parameters in the scattering problem mainly shift and distort the interference fringe pattern.
[…]
Several promising directions for extending the present work naturally emerge. First, investigating higher-spin fields or nonminimally coupled scalar fields in the Frolov spacetime would help assess the universality of the patterns identified herein. Second, exploring rotating generalizations of regular black holes could uncover qualitatively new signatures in absorption and scattering, as the interplay between regular cores and frame-dragging effects may introduce novel physical behaviors. Such extensions would further deepen our understanding of the interaction between waves and regular black holes, providing valuable insights for future observational tests of gravitational theories. Beyond the present static setup, it is natural to consider charged/massive scalar probes in charged backgrounds; see, e.g., Ref. [32] and Ref. [33]. It would also be interesting to connect black hole scattering/absorption calculations to wave-optics gravitational-wave lensing, where spin effects can enter naturally [35].
ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China (NSFC) under Grant nos. 12235019, 12275106 and Shandong Provincial Natural Science Foundation under grant No. ZR2024QA032.